OpenStudy (amistre64):

so what does the "a" contribute in a taylor series? P(x) = f(a)+f'(a)(x-a)+f''(a)(x-a)^2 ...

OpenStudy (lalaly):

its the x0 ... when they say find the taylor polynomial about a point x0

OpenStudy (amistre64):

right .... they call it a basis point in some texts; and when it modifies the x part its translates into a shift in the graph ... I was curled up with cozy calculus book last night and finally realized what a taylor series was. Before that it was rather enigmatic to me. But I was curious about the x-a parts ...

OpenStudy (lalaly):

can i ask yu a question now?...(A)

OpenStudy (amistre64):

I made a cos(x) polynomial with a=0 .. but was curious about what would happen if I made a= something else but feel asleep and dreamed of rabid gators or some such :)

OpenStudy (amistre64):

you can ask

OpenStudy (lalaly):

are you good at numerical analysis?

OpenStudy (amistre64):

i dont know; i might be, but I never sat down and tried to determine what numerical analysis actually is :)

OpenStudy (lalaly):

and btw when a=0 its calld a maclaurin series

OpenStudy (lalaly):

ok i'll show u the question,

OpenStudy (amistre64):

nice :) I found out last night that the poly expressed the degrees in radians tho ... would our a=the radian version of 42 then?

OpenStudy (lalaly):

yess which is 7pi/30

OpenStudy (amistre64):

and would we be polying up this with the derivatives of cos(pi/4) ??

OpenStudy (anonymous):

sorry to be rude ha but amistre i still need help man could you help when u done here

OpenStudy (amistre64):

cos(a) = 1/sqrt(2) -sin(a) = -1/sqrt(2) -cos(a) = 1/sqrt(2) ... like that?

OpenStudy (amistre64):

ill take a look at it when i can :)

OpenStudy (lalaly):

im confused :( :'(

OpenStudy (amistre64):

cos(d-a) then.... that doesnt seem like an easy thing to check unless you already know the sin and cos amounts for (7pi/30-pi/4) ...

OpenStudy (amistre64):

28/120 30/120 are pretty close, so id assume its evalulated at pi/4

OpenStudy (lalaly):

i hav the solution but its confusing me, im stuck

OpenStudy (amistre64):

lets use a=pi/4 (30pi/120) to approximate a poly for cos(7pi/30), (28pi/120) ok?

OpenStudy (lalaly):

okaaay

OpenStudy (amistre64):

lets take a generic poly set up and derive it down then we can equate it to the derivaites of cos(45): \[P(x) = ax^8+bx^7+cx^6+dx^5+ex^4+fx^3+gx^2+hx+i\] \[P1(x) = 8ax^7+7bx^6+6cx^5+5dx^4+4ex^3+3fx^2+2gx+h\] \[P_1(x) = 8ax^7+7bx^6+6cx^5+5dx^4+4ex^3+3fx^2+2gx+h\] \[P_2(x) = 56ax^6+42bx^5+30cx^4+20dx^3+12ex^2+6fx+2g\] \[P_3(x) = 336ax^5+210bx^4+120cx^3+60dx^2+24ex+6fx\] etc right?

OpenStudy (amistre64):

eventually we get down to factorials; since each coefficient is the product of each successive deriving: 8.7.6.5.4.3.2.1 for the x^8 part right?

OpenStudy (lalaly):

thats what i got first i find the exact value of cos 42 7 digits then i use taylor poly till thers a difference a number in 10^-6 then cos 42 - P3(x) = error and thats it :S

OpenStudy (amistre64):

0. 1/sqrt(2) 1. -1/sqrt(2) 2. -1/sqrt(2) 3. 1/sqrt(2) 4. 1/sqrt(2) 5. -1/sqrt(2) 6. -1/sqrt(2) 7. 1/sqrt(2) 8. 1/sqrt(2) \[ [cos(45)]_8=P_8=8!a=1/\sqrt{2};\ a=1/8!\sqrt{2}\] \[ [cos(45)]_7=P_7=\frac{8!}{8!/\sqrt{2}}+7!b=1/\sqrt{2}\] \(=\sqrt{2}+7!b=\cfrac{1}{\sqrt{2}}\) \( 7!b=\cfrac{1}{\sqrt{2}}-\sqrt{2}\) \( b=\left(\cfrac{1}{\sqrt{2}}-\sqrt{2}\right)/7! \) that gets messy real quick like it seems :)

OpenStudy (amistre64):

and typing it in latex format is a pain...

OpenStudy (amistre64):

at any rate; that is where we get the factorial denominators from P(x)=\(\frac{}{0!}+\frac{}{1!}x+\frac{}{2!}x^2+\frac{}{3!}x^3+\frac{}{4!}x^4+\frac{}{5!}x^5+\frac{}{6!}x^6+\frac{}{7!}x^7+\frac{}{8!}x^8\) just curious; but is this when we use the (x-a)^# stuff?

OpenStudy (lalaly):

yesss

OpenStudy (amistre64):

and the tops of the fractions here equal the derivative values right? f(a), f'(a), f''(a), ... correct?

OpenStudy (lalaly):

yess correct

OpenStudy (amistre64):

\[P(x)=cos(a)-\frac{sin(a)}{1!}(x-\frac{\pi}{4})-\frac{cos(a)}{2!}(x-\frac{\pi}{4})^2+\frac{sin(a)}{3!}(x-\frac{\pi}{4})^3...\]

OpenStudy (amistre64):

\[P(x)=\frac{1}{\sqrt{2}}-\frac{(x-\frac{\pi}{4})}{\sqrt{2}}-\frac{(x-\frac{\pi}{4})^2}{2!\sqrt{2}}+\frac{(x-\frac{\pi}{4})^3}{3!\sqrt{2}}+\frac{(x-\frac{\pi}{4})^4}{4!\sqrt{2}}-\frac{(x-\frac{\pi}{4})^5}{5!\sqrt{2}}\]

OpenStudy (amistre64):

add more poly terms as needed :)

OpenStudy (anonymous):

In a power series we talk about a power series being "centered" at a given point. In many cases the power series won't converge absolutely on the entire domain, it will only converge within a "radius of convergence" and it does so centered at "a".

OpenStudy (amistre64):

id go to 8 and test it out; but that many just goes off the screen

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Radius_of_convergence

OpenStudy (amistre64):

or stop at 5 and test it out :)

OpenStudy (amistre64):

so the "a" is a way to move our poly to the desired position and fit it to our curve right?

OpenStudy (anonymous):

Except in some cases the radius of convergence will be infinite and so it doesnt matter what "a" you choose.

OpenStudy (amistre64):

im gonna have to plug this into me calculator to assess its veracity :)

OpenStudy (anonymous):

the function will be "centered" around a im just saying it will converge everywhere if the radius of convergence is infinite.

OpenStudy (anonymous):

you would need to graph it to check this.

OpenStudy (amistre64):

my graphs dont line up when I do the x-a parts .....

OpenStudy (anonymous):

try it for sin(x) then choose any a you like

OpenStudy (anonymous):

\[\sum _{n=0}^{\infty }{\frac { sin^{ \left( n \right) } \left( 10 \right) \left( x-10 \right) ^{n}}{n!}} = sin(x)\]

OpenStudy (amistre64):

when I use the cos(0) for my derivative bits I get it spot on to 6 decimals

OpenStudy (anonymous):

I just chose a = 10 but it doesn't matter

OpenStudy (anonymous):

the - a in the factor (x-a) just offsets the "centering" in the choice of coefficients. without it, it would be sin(x + a)

OpenStudy (anonymous):

Anyways the trig functions sin, and cos have infinite radiuses of convergence.

OpenStudy (anonymous):

Thats why "a" for those functions don't matter

OpenStudy (amistre64):

ohhh.....